In 6 years, Sylvia will be 6/5 Hue's age. Four years ago, Sylvia was 10 years less than twice Hue's age. How old are Sylvia and Hue now?
Answer by TimothyLamb(4379)

x = sylvia's age now
y = hue's age now
---
x + 6 = (6/5)y
x - 4 = 2(y - 4) - 10
---
x + 6 = (6/5)y
x - 4 = 2y - 8 - 10
---
put the system of linear equations into standard form
---
x - 1.2y = 6
x - 2y = -14
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = sylvia's age now = 36
y = hue's age now = 25
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php